OpenStudy (anonymous):

which of the following is equivalent to.... 1-cos^2 ϴ/cos^2 ϴ? a.sec^2ϴ b.(csc^2ϴ)-1 c.tan^2 ϴ d.sin^2ϴ answer is C. but how?

OpenStudy (anonymous):

is it (1-cos^2x)/cos^2x or without brackets?

OpenStudy (cwrw238):

if its with brackets (1-cos^2x)/cos^2x now (1-cos^2x) = sin^2x so the above = sin^2x / cos^2x = tan^2x C

OpenStudy (anonymous):

its without brackets.... i know the answer but i would like to know how??

OpenStudy (anonymous):

without brackets doesn't make sense\[1-\cos ^2\theta/\cos ^2\theta=1-1=0\]

OpenStudy (anonymous):

\[1-\cos^2\theta/\cos^2\theta \]

OpenStudy (anonymous):

that's exactly how it is on the paper without brackets..

OpenStudy (cwrw238):

well the guys who wrote the paper must have forgotten the brackets

OpenStudy (cwrw238):

Jithinkv is correct